A reactor, which will contain corrosive liquid, has been design. If the reactor is made of mild steel, the initial installed cost will be $6000, and the useful life period will be 4 years. Since stainless steel is highly resistant to the corrosive action of the liquids, stainless steel, as the material of construction, has been proposed as an alternative to mild steel. the stainless steel reactor would have an initial installed cost of $16000. The scrape value at the end off the useful life would be zero for either type of reactor, and both could be replaced at a cost equal to the original price. On the basis of equal capitalized cost for both types of reactors, what should be the useful life period for the stainless steel reactor if money is worth 4 percent compounded annually?
1- The capitalized cost
2-the useful life period for the stainless steel reactor

i need a solution within half an hour
Answer #1

Son Capitalized cost = Cu + Co-Cs (1+i)., given C, = $6000 Cs= $0, n=4 a is 0.04 fuy. C.C - 6000 + 6000 (104)-, (.6= $4132 3